Please use this description to calculate the values requested below (correct to 3 decimal places)...

The athletic departments at 10 randomly selected U.S. universities were asked by the Equal Employment Opportunity Commission to state what percentage of their nursing scholarships were presently held by women. The responses were 5, 4, 2, 1, 1, 2, 10, 2, 3, 5.

Geometric mean

Midhinge Trimmed mean (hint: use Minitab)

Standard deviation

IQR

Variance

MAD

CV (hint: write this as a percentage without the % symbol, not a proportion)

Skewness

Kurtosis

p-value for the Anderson-Darling normality test

Upper outer fence of the boxplot

Which distribution is least likely to be skewed to the right by high values?
Answer #1

Mean (X bar) = Sum of Values /n

and

Hence, from the given data,

Mean = 3.5

Std Dev (s) = 2.718

Variance = s2 = 7.389

Median = 2.5

Quartile 1 (Q1) = 2

Quartile 3 (Q3) = 4.75

Midhinge Trimmed mean = (Q3+Q1)/2 = 3.375

IQR = Q3 – Q1 = 2

Skewness = 1.660

Kurtosis = 3.211